# QCD phase transition at finite temperature and chemical potential with the non-extensive statistics

### v1

@@ III.2 Finite chemical potential
For the system with finite chemical potential, we only consider the QCD phase transition in the non-chiral limit case. The baryon, strangeness and electric charge potentials are, respectively, set as $\mu_{B}=$ 398.2 MeV, $\mu_{S}=$ 89.5 MeV and $\mu_{Q}=$ 0 MeV, which correspond to the chemical potentials of the system created in the central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7 GeV [65]. Figure 4 shows the temperature dependence of $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $\Delta_{q}$, $(C_{V})_{q}/T^{3}$ and $(c_{s}^{2})_{q}$. The solid and dashed black (blue) lines correspond to $q=$ 1.05 and $q=$ 1 with zero (finite) chemical potential, respectively. In the figure, $T_{c}$ is taken to be 156.5 MeV [64], $\it\Gamma$ is set as 0.05 $T_{c}$. The qualitative behaviors of the thermodynamic quantities are summarized as follows.
(i) $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $(C_{V})_{q}/T^{3}$ and $\Delta_{q}$ are sensitive to the deviation of the chemical potential from zero and they become large both in the hadronic and QGP phases with the increase of $\mu$. This is due to the reason that at a given temperature a larger population is expected for the system with finite chemical potential. A similar increase was observed for the normalized pressure and trace anomaly obtained from the Dyson-Schwinger equations when the values of quark chemical potential were increased [66].
(ii) With the increase of the chemical potential, the sound velocity squared $(c_{s}^{2})_{q}$ increases in the vicinity of $T_{c}$ and decreases at lower temperature, which is different from the chemical potential dependence of $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $(C_{V})_{q}/T^{3}$ and $\Delta_{q}$.
@@
Figure 4: $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $\Delta_{q}$, $(C_{V})_{q}/T^{3}$ and $(c_{s}^{2})_{q}$ as a function of $T/T_{c}$ in the non-chiral limit case. The solid and dashed black lines correspond to $q=$ 1.05 and $q=$ 1 with zero chemical potential, respectively. The solid and dashed blue lines correspond to $q=$ 1.05 and $q=$ 1 with $\mu_{B}=$ 398.2 MeV, $\mu_{S}=$ 89.5 MeV and $\mu_{Q}=$ 0 MeV, respectively.
@@ III.3 Comparison with other models
Figure 5: Upper panel: $P_{q}/T^{4}$ as a function of $T$ with zero chemical potential for $q=$ 1.05 (the solid line) and 1 (the dashed line) in the non-chiral limit case. The results from the HRG model, the LQCD and the Thermal-Fist fit are, respectively, shown as the dotted line, the solid black circles and the empty markers. Lower panel: $P_{q}/T^{4}$ as a function of $T$ with finite chemical potential for $q=$ 1.05 (the solid line) and $q=$ 1 (the dashed line) in the non-chiral limit case. The results from the HRG model and the Thermal-Fist fit are, respectively, shown as the dotted line and the empty markers.
@@
The upper panel of Fig. 5 displays the comparison of the temperature dependence of $P_{q}/T^{4}$ calculated with zero chemical potential for different models. Our results with $q=$ 1.05 and $q=$ 1 in the non-chiral limit case are, respectively, presented as the solid and dashed lines. The result from the HRG model is shown as the dotted line, while the result from the LQCD is presented as the solid black circles. The dependence extracted from the fit of experimental yields of hadrons at the central centrality in pp collisions at $\sqrt{s}=$ 7 TeV [67, 68, 69] and 13 TeV [70, 71, 72], p-Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV [73, 74, 75], Pb-Pb collisions at $\sqrt{s_{\rm NN}}=$ 2.76 TeV [76, 77, 78, 79] and 5.02 TeV [80, 81], and Xe-Xe collisions at $\sqrt{s_{\rm NN}}=$ 5.44 TeV [82] using the Thermal-Fist package [58] are shown as empty markers. Thermal-Fist is a C++ package designed for the analyses of particle productions in relativistic heavy-ion collisions within the framework of the HRG model. In the package, there are several options regarding the ensemble with which to treat the baryon number, strangeness and electric charge. In this work, we choose the grand canonical ensemble, which is widely applied to heavy-ion collisions [83, 84, 85, 86]. It is shown that our result with $q=1$ agrees well with the HRG model and the LQCD data up to $T\sim$ 150 MeV. In our model, $P_{q}(T)$ quickly approaches the pressure at the hadronic phase with $T<T_{c}-\it\Gamma$. When $q\rightarrow 1$, the Tsallis statistics approaches the BG statistics valid for systems under the thermal equilibrium. Thus, the agreement at low temperature is reasonably expected as the HRG model and the LQCD calculations are based on equilibrium considerations. Moreover, our result with $q=1$ is consistent with the results from the Thermal-Fist fit to the hadron yields at the central centrality in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=$ 2.76 TeV (triangle-left) and 5.02 TeV (triangle-up), and in Xe-Xe collisions at $\sqrt{s_{\rm NN}}=$ 5.44 TeV (triangle-down). However, it overestimates the results from the Thermal-Fist fit to the hadron yields in p-Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV (triangle-right), and in pp collisions at $\sqrt{s}=$ 7 TeV (empty circle) and 13 TeV (empty square). In pp and p-Pb collisions, the temperature of the system is around 160 MeV, which is in the phase transition region $\lvert T-T_{c}\rvert<\it\Gamma$, leading to the overestimate of our model for the pressure. At high temperature, our result approaches the SB limit of the pressure for idea parton gas. However, in LQCD, partons interact with each other, which leads to the deviation from the limit. In the HRG model, only the hadron gas is considered and the phase transition is absent, thus the pressure approaches infinity at high temperature.
Table 1: Chemical potentials in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7, 11.5 and 200 GeV.
@@
The lower panel of Fig. 5 shows the comparison of the temperature dependence of $P_{q}/T^{4}$ calculated with three sets of finite chemical potentials (see table 1) for different models. They correspond to the chemical potentials in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7, 11.5 and 200 GeV [65]. Our results with $q=$ 1.05 and $q=$ 1 in the non-chiral limit case are, respectively, presented as the solid and dashed lines. The results from the HRG model are shown as the dotted line. The dependence extracted from the Thermal-Fist fit of experimental yields of hadrons in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7, 11.5 and 200 GeV [65, 87, 88, 89] are shown as empty blue, red and black markers, respectively. It is found that our results with $q=1$ agree well with the HRG model up to $T\sim$ 150 MeV. Moreover, our results with $q=1$ are consistent with the results from the Thermal-Fist fit to the hadron yields in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7 GeV (triangle-right) and 11.5 GeV (triangle-left). However, they overestimate the result from the Thermal-Fist fit to the hadron yields in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV (triangle-up). The temperature of the system in Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV is around 160 MeV, falling in the phase transition region and leading to this overestimate.
@@ IV Conclusions
In this work, by applying the non-extensive correction to the EoS in the parton (hadron resonance) gas at high (low) temperature and making a smooth interpolation of these two equation of states, we have constructed a non-extensive version of EoS to investigate the properties of QCD phase transition with zero chemical potential in the chiral and non-chiral limit cases. We have also investigated the QCD phase transition with finite chemical potential in the non-chiral limit case. We find that the dimensionless quantities $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $(C_{V})_{q}/T^{3}$ and $\Delta_{q}$ are sensitive to the deviation of $q$ from unity and they become large both in the hadronic and QGP phases with the increase of $q$. In the chiral limit case, $(c_{s}^{2})_{q}$ is insensitive to the deviation of $q$ from unity. However, in the non-chiral limit case, this deviation leads to nontrivial corrections of $(c_{s}^{2})_{q}$ in the vicinity of the critical point and at lower temperatures. Moreover, these dimensionless thermodynamic quantities are sensitive to the deviation of $\mu$ from zero and they become large both in the hadronic and QGP phases with the increase of $\mu$. However, $(c_{s}^{2})_{q}$ increases in the vicinity of the phase transition temperature and decreases at lower temperature. Finally, in the non-chiral limit case, we observe that our results with $q=1$ agree well with the LQCD, the HRG and the results form the Thermal-Fist fit to the hadron yield in high energy nuclear collisions in the low temperature region up to $T\sim 150$ MeV. These findings will deepen our understanding for the properties of the QCD phase transition at finite temperature and chemical potential both in the extensive and non-extensive statistics.

### v2

@@ III.2 Finite chemical potential
The QCD phase transition is also investigated at finite chemical potential. The baryon, strangeness, and electric charge potentials are, respectively, set as $\mu_{B}=$ 398.2 MeV, $\mu_{S}=$ 89.5 MeV, and $\mu_{Q}=$ 0 MeV, which correspond to the chemical potentials of the system created in the central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7 GeV [71]. In each subplot of Fig. 1, the solid and dashed blue curves, respectively, show the temperature dependence of the thermodynamic quantities with finite chemical potential for $q=$ 1.05 and $q=$ 1. For the case with finite chemical potential, $T_{c}$ and $\it\Gamma$ are set to be the same values as those for the case with vanishing chemical potential. The qualitative behaviors of the thermodynamic quantities are summarized as follows.
(i) $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $(C_{V})_{q}/T^{3}$, and $\Delta_{q}$ are sensitive to the deviation of the chemical potential from zero and they become large both in the hadronic and QGP phases with the increase of $\mu$. This is due to the reason that at a given temperature a larger population is expected for the system with finite chemical potential. A similar increase was observed for the normalized pressure and trace anomaly obtained from the Dyson-Schwinger equations when the values of quark chemical potential were increased [72].
(ii) With the increase of the chemical potential, the squared speed of sound $(c_{s}^{2})_{q}$ increases in the vicinity of $T_{c}$ and decreases at lower temperatures, which is different from the chemical potential dependence of $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $(C_{V})_{q}/T^{3}$ and $\Delta_{q}$.
@@ III.3 Comparison with other models
Figure 3: Upper panel: $P_{q}/T^{4}$ as a function of $T$ with zero chemical potential for $q=$ 1.05 (the solid line) and 1 (the dashed line). The results from the HRG model, the LQCD, and the Thermal-Fist fit are, respectively, shown as the dotted line, the solid black circles, and the empty markers. The inset is the temperature dependence of $P_{q}/T^{4}$ in the region with 136 MeV <T<174<T<174 MeV. Lower panel: $P_{q}/T^{4}$ as a function of $T$ with finite chemical potential for $q=$ 1.05 (the solid line) and $q=$ 1 (the dashed line). The results from the HRG model and the Thermal-Fist fit are, respectively, shown as the dotted line and the empty markers.
@@
The upper panel of Fig. 3 displays the comparison of the temperature dependence of $P_{q}/T^{4}$ calculated with zero chemical potential for different models. Our results with $q=$ 1.05 and $q=$ 1 are, respectively, presented as the solid and dashed lines. The result from the HRG model is shown as the dotted line, while the result from the LQCD is presented as the solid black circles. The dependence extracted from the fit of experimental yields of hadrons at the central centrality in pp collisions at $\sqrt{s}=$ 7 TeV [73, 74, 75] and 13 TeV [76, 77, 78], p-Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV [79, 80, 81], Pb-Pb collisions at $\sqrt{s_{\rm NN}}=$ 2.76 TeV [82, 83, 84, 85] and 5.02 TeV [86, 87], and Xe-Xe collisions at $\sqrt{s_{\rm NN}}=$ 5.44 TeV [88] using the Thermal-Fist package [63] are shown as empty markers. Thermal-Fist is a C++ package designed for the analyses of particle productions in relativistic heavy-ion collisions within the framework of the HRG model. In the package, there are several options regarding the ensemble with which to treat the baryon number, strangeness, and electric charge. In this work, we choose the grand canonical ensemble, which is widely applied to heavy-ion collisions [89, 90, 91, 92]. It is shown that our result with $q=1$ agrees well with the HRG model and the LQCD data up to $T\sim$ 150 MeV. In our model, $P_{q}(T)$ quickly approaches the pressure at the hadronic phase with $T<T_{c}-\it\Gamma$. When $q\rightarrow 1$, the Tsallis statistics approaches the BG statistics valid for systems under the thermal equilibrium. Thus, the agreement at low temperature is reasonably expected as the HRG model and the LQCD calculations are based on equilibrium considerations. Moreover, our result with $q=1$ is consistent with the results from the Thermal-Fist fit to the hadron yields at the central centrality in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=$ 2.76 TeV (triangle-left) and 5.02 TeV (triangle-up), and in Xe-Xe collisions at $\sqrt{s_{\rm NN}}=$ 5.44 TeV (triangle-down). However, it obviously overestimates the results from the Thermal-Fist fit to the hadron yields in p-Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV (triangle-right), and in pp collisions at $\sqrt{s}=$ 7 TeV (empty circle) and 13 TeV (empty square). In pp and p-Pb collisions, the temperature of the system is around 160 MeV, which is in the phase transition region $\lvert T-T_{c}\rvert<\it\Gamma$, leading to the overestimate of our model for the pressure. At high temperature, our result approaches the SB limit of the pressure for idea parton gas. However, in LQCD, partons interact with each other, which leads to the deviation from the limit. In the HRG model, only the hadron gas is considered and the phase transition is absent, thus the pressure approaches infinity at high temperature.
Table 1: Chemical potentials in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7, 11.5, and 200 GeV.
@@
The lower panel of Fig. 3 shows the comparison of the temperature dependence of $P_{q}/T^{4}$ calculated with three sets of finite chemical potentials (see table 1) for different models. They correspond to the chemical potentials in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7, 11.5, and 200 GeV [71]. Our results with $q=$ 1.05 and $q=$ 1 are, respectively, presented as the solid and dashed lines. The results from the HRG model are shown as the dotted line. The dependence extracted from the Thermal-Fist fit of experimental yields of hadrons in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7, 11.5, and 200 GeV [71, 93, 94, 95] are shown as empty blue, red, and black markers, respectively. It is found that our results with $q=1$ agree well with the HRG model up to $T\sim$ 150 MeV. Moreover, our results with $q=1$ are consistent with the results from the Thermal-Fist fit to the hadron yields in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7 GeV (triangle-right) and 11.5 GeV (triangle-left). However, they overestimate the result from the Thermal-Fist fit to the hadron yields in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV (triangle-up). The temperature of the system in Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV is around 160 MeV, falling in the phase transition region and leading to this overestimate.
@@ IV Conclusions
In this work, by applying the non-extensive correction to the EoS in the parton (hadron resonance) gas at high (low) temperature and making a smooth interpolation of these two equation of states, we have constructed a non-extensive version of EoS to investigate the properties of QCD phase transition with both zero and non-zero chemical potentials. We find that the dimensionless quantities such as $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $(C_{V})_{q}/T^{3}$, and $\Delta_{q}$ are highly sensitive to the non-extensivity parameter $q$. As $q$ deviates from unity, these quantities increase in both the hadronic and QGP phases. Furthermore, this deviation introduces nontrivial corrections to $(c_{s}^{2})_{q}$ near the critical point and at lower temperatures. These thermodynamic quantities also exhibit significant sensitivity to the chemical potential $\mu$. They increase with $\mu$ in both phases. Specifically, $(c_{s}^{2})_{q}$ rises near the phase transition temperature but decreases at lower temperatures. Finally, we observe that our results with $q=1$ agree well with the LQCD, the HRG, and the results form the Thermal-Fist fit to the hadron yield in high energy nuclear collisions in the low temperature region up to $T\sim 150$ MeV. These findings will deepen our understanding for the properties of the QCD phase transition at finite temperature and chemical potential both in the extensive and non-extensive statistics.